Which of the following are solutions to the quadratic equation? Check all that apply. x2+2x-8-0 nB. 28
Answer #1

Solution: Option   A\: \: \: ,\: \: \: E

x^{2}+2x-8=0

2^{2}+2\times 2-8=4+4-8=8+8=0

\therefore 2   is a solution

28^{2}+2\times 28-8\neq

\therefore 28   is not a solution

1^{2}+2\times 1-8=1+2-8=-5\neq 0

\therefore 1   is not a solution

\left ( -1 \right )^{2}+2\times \left ( -1 \right )-8=1-2-8=-9\neq 0

\therefore -1   is not a solution

\left ( -4 \right )^{2}+2\times \left ( -4 \right )-8=16-8-8=0

\therefore -4   is a solution

4^{2}+2\times 4-8=16

\therefore 4   is not a solution